OpenStudy (anonymous):

Let ρAl represent the density of aluminum and ρFe that of iron. Find the radius of a solid aluminum sphere that balances a solid iron sphere of radius rFe on an equal-arm balance. (Use any variable or symbol stated above as necessary.)

OpenStudy (ash2326):

ρAl ρFe let the radius of aluminimu= rAl radius of iron = rFe Volume of iron sphere \[ Viron= 4/3 {\pi} {rFe}^3\] \[ Valu= 4/3 \pi {rAl}^3\] we now mass is geiven as= density * volume Miron= Malum so \[ ρFe * 4/3 \pi {rfe}^3= ρAl * 4/3 \pi {rAl}^3\] or \[ rAl = rfe \sqrt[3] { ρFe/ρAl }\]